# A Thermodynamically Consistent Free Boundary Model for Two-Phase Flows in an Evolving Domain with Moving Contact Lines, Variable Contact Angles

### v1

@@ 2.3   The role of dynamic boundary conditions in our model
  • For $K=0$ and $L\in(0,\infty)$, the non-convective bulk-surface Cahn–Hilliard equation was proposed in [59] to interpolate between the models associated with $K=0$, $L=0$ and ${K=0}$, ${L=\infty}$. In this case, a transfer of material between bulk and surface will occur, and the number $L^{-1}$ is related to the rate of absorption (or adsorption) of bulk material by the boundary (cf. [59]).
@@ 2.4   Variants of our model with a static boundary
$\partial_{t}\big{(}\rho(\phi){\mathbf{u}}\big{)}+\textnormal{div}\,\big{(}{\mathbf{u}}\otimes(\rho(\phi){\mathbf{u}}+\mathbf{J})\big{)}=\textnormal{div}\,(\mathbf{T}),\quad\textnormal{div}\,{\mathbf{u}}=0,$ in $\quad\text{in }Q_{T},$ (2.4a)
$\partial_{t}\phi+\textnormal{div}\,(\phi{\mathbf{u}})=\textnormal{div}\,(m_{\Omega}(\phi)\nabla\mu),$ in $\quad\text{in }Q_{T},$ (2.4b)
$\mu=-\varepsilon\Delta\phi+\varepsilon^{-1}F^{\prime}(\phi),$ in $\quad\text{in }Q_{T},$ (2.4c)
$\partial_{t}\psi+\textnormal{div}_{\Gamma}(\psi{\mathbf{u}}_{\tau})=\textnormal{div}_{\Gamma}(m_{\Gamma}(\psi)\nabla_{\Gamma}\theta)-\beta m_{\Omega}(\phi)\partial_{\mathbf{n}}\mu,$ on $\quad\text{on }S_{T},$ (2.4d)
$\theta=-\delta\Delta_{\Gamma}\psi+\delta^{-1}G^{\prime}(\psi)+\alpha\varepsilon\partial_{\mathbf{n}}\phi,$ on $\quad\text{on }S_{T},$ (2.4e)
$K\varepsilon\partial_{\mathbf{n}}\phi=\alpha\psi-\phi,\quad Lm_{\Omega}(\phi)\partial_{\mathbf{n}}\mu=\beta\theta-\mu,\quad K,L\in[0,\infty],$ on $\quad\text{on }S_{T},$ (2.4f)
$\big{[}\mathbf{S}{\mathbf{n}}+\nabla_{\Gamma}q+\gamma_{\tau}{\mathbf{u}}\big{]}_{\tau}=\Big{[}\theta\nabla_{\Gamma}\psi+\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\,{\mathbf{u}}\Big{]}_{\tau},$ on $\quad\text{on }S_{T},$ (2.4g)
$\textnormal{div}_{\Gamma}{\mathbf{u}}=0,\quad{\mathbf{u}}\cdot{\mathbf{n}}=0,$ on $\quad\text{on }S_{T},$ (2.4h)
@@ 3   Model derivation
where for any $(t,{\mathbf{x}})\in\overline{Q_{T}}$, the flow map $[0,T]\ni s\mapsto\eta_{s}(t,{\mathbf{x}})\in\overline{\Omega(s)}$ is the unique solution of the ordinary differential equation
@@
[…]
@@ 3.1   Considerations based on local mass balance laws
Here, $\widehat{\mathbf{J}}_{i}$ are the mass fluxes corresponding to the motion of the two materials in the bulk. In our model, we also want to account for a transfer of material between bulk and surface. This is, for instance, needed in order to describe absorption processes. Therefore, we also consider functions $\sigma_{i}:S_{T}\to\mathbb{R}$, $i=1,2$, which represent the mass densities of the substrates on the free boundary. In general, the materials on the surface might differ from those in the bulk. This is the case, for example, if the materials are transformed by chemical reactions occurring at the boundary. Therefore, we interpret the densities $\sigma_{i}$ as independent functions that are not necessarily the traces of the densities $\rho_{i}$, respectively. We assume that the the following mass balance law holds on the free boundary:
@@
$\rho(\partial_{t}{\mathbf{u}}+{\mathbf{u}}\cdot\nabla{\mathbf{u}})&=\textnormal{div}\,\,\tilde{\mathbf{T}}+(\textnormal{div}\,\,\mathbf{J}){\mathbf{u}}=\textnormal{div}\,(\tilde{\mathbf{T}}+{\mathbf{u}}\otimes\mathbf{J})-\nabla{\mathbf{u}}\;\mathbf{J}\quad\text{in $Q_{T}$.}\end{aligned}$ (3.23)
@@
$\textnormal{div}\,{\mathbf{u}}=\textnormal{div}\,\left(\frac{\rho_{1}}{\tilde{\rho}_{1}}{\mathbf{u}}_{1}+\frac{\rho_{2}}{\tilde{\rho}_{2}}{\mathbf{u}}_{2}\right)=-\partial_{t}\left(\frac{\rho_{1}}{\tilde{\rho}_{1}}+\frac{\rho_{2}}{\tilde{\rho}_{2}}\right)=-\partial_{t}1=0&&\quad\text{in $Q_{T}$,}\end{aligned}$ (3.26)
@@
$&\textnormal{div}_{\Gamma}\,{\mathbf{u}}=\textnormal{div}_{\Gamma}\left(\frac{\sigma_{1}}{\tilde{\sigma}_{1}}{\mathbf{u}}_{1}+\frac{\sigma_{2}}{\tilde{\sigma}_{2}}{\mathbf{u}}_{2}\right)\\ &\quad=-\partial_{t}^{\circ}\left(\frac{\sigma_{1}}{\tilde{\sigma}_{1}}+\frac{\sigma_{2}}{\tilde{\sigma}_{2}}\right)+{\mathbf{u}}_{\tau}\cdot\nabla_{\Gamma}\left(\frac{\sigma_{1}}{\tilde{\sigma}_{1}}+\frac{\sigma_{2}}{\tilde{\sigma}_{2}}\right)+\left(\frac{\widehat{\mathbf{J}}_{\Gamma,1}}{\tilde{\sigma}_{1}}+\frac{\widehat{\mathbf{J}}_{\Gamma,2}}{\tilde{\sigma}_{2}}\right)\cdot{\mathbf{n}}\\ &\quad=\left(\frac{\widehat{\mathbf{J}}_{\Gamma,1}}{\tilde{\sigma}_{1}}+\frac{\widehat{\mathbf{J}}_{\Gamma,2}}{\tilde{\sigma}_{2}}\right)\cdot{\mathbf{n}}=0&&\quad\text{on $S_{T}$.}\end{aligned}$ (3.27)
@@
$\tilde{\mathbf{J}}_{\Gamma,i}$ $=\frac{\tilde{\mathbf{J}}_{\Gamma,i}}{\tilde{\sigma}_{i}},$ $\quad i=1,2,$ $\qquad\mathbf{J}_{\psi}$ $=\tilde{\mathbf{J}}_{\Gamma,2}-\tilde{\mathbf{J}}_{\Gamma,1}$ on $S_{T}$, (3.29)
@@
$\textnormal{div}\,\,\mathbf{J}&=\frac{1}{2}\bigg{[}\tilde{\rho}_{2}\textnormal{div}\,\,\tilde{\mathbf{J}}_{2}+\tilde{\rho}_{1}\textnormal{div}\,\,\tilde{\mathbf{J}}_{1}\bigg{]}+\frac{1}{2}\bigg{[}\tilde{\rho}_{2}\textnormal{div}\,\,\tilde{\mathbf{J}}_{2}+\tilde{\rho}_{1}\textnormal{div}\,\,\tilde{\mathbf{J}}_{1}\bigg{]}\\[4.30554pt] &=\frac{\tilde{\rho}_{2}-\tilde{\rho}_{1}}{2}\textnormal{div}\,\,\tilde{\mathbf{J}}_{2}-\frac{\tilde{\rho}_{2}-\tilde{\rho}_{1}}{2}\textnormal{div}\,\,\tilde{\mathbf{J}}_{1}=\frac{\tilde{\rho}_{2}-\tilde{\rho}_{1}}{2}\textnormal{div}\,\,\mathbf{J}_{\phi}&&\quad\text{in $Q_{T}$.}\end{aligned}$ (3.35)
@@
Assuming that $\phi$ and $\psi$ only attain values in the interval $[-1,1]$, we use (3.9), (3.10) and (3.17) to conclude that the density $\rho=\rho(\phi)$ is given by
@@ 3.2   Local energy dissipation laws
[…]
@@
$0&\geq\frac{\mathrm{d}}{\mathrm{d}t}\left[\int_{V(t)}e_{\Omega}({\mathbf{u}},\phi,\nabla\phi)\;\mathrm{d}{\mathbf{x}}+\int_{\partial V(t)\cap\Gamma(t)}e_{\Gamma}(\phi,\psi,\nabla_{\Gamma}\psi)\;\mathrm{d}\mathcal{H}^{d-1}\right]\\ &\qquad+\int_{V(t)}\textnormal{div}\,\,\mathbf{J}_{e}\;\mathrm{d}{\mathbf{x}}+\int_{\partial V(t)\cap\Gamma(t)}\textnormal{div}_{\Gamma}\,\mathbf{K}_{e}+H\mathbf{K}_{e}\cdot{\mathbf{n}}-\mathbf{J}_{e}\cdot{\mathbf{n}}\;\mathrm{d}\mathcal{H}^{d-1}\,.\end{aligned}$ (3.44)
@@
$0&\geq\int_{V(t)}\partial_{t}e_{\Omega}({\mathbf{u}},\phi,\nabla\phi)+\textnormal{div}\,\big{(}e_{\Omega}({\mathbf{u}},\phi,\nabla\phi){\mathbf{u}}\big{)}+\textnormal{div}\,\,\mathbf{J}_{e}\;\mathrm{d}{\mathbf{x}}\\ &\quad+\int_{\partial V(t)\cap\Gamma(t)}\partial_{t}^{\circ}e_{\Gamma}(\phi,\psi,\nabla_{\Gamma}\psi)\;\mathrm{d}\mathcal{H}^{d-1}\\ &\quad+\int_{\partial V(t)\cap\Gamma(t)}\textnormal{div}_{\Gamma}\,\mathbf{K}_{e}+H\mathbf{K}_{e}\cdot{\mathbf{n}}-\mathbf{J}_{e}\cdot{\mathbf{n}}\;\mathrm{d}\mathcal{H}^{d-1}.\end{aligned}$ (3.47)
@@ 3.3.1 Computations in the bulk
$0&\geq\textnormal{div}\,\big{[}\mathbf{J}_{e}-\tfrac{1}{2}\left|{\mathbf{u}}\right|^{2}\mathbf{J}+\mathbf{T}^{T}{\mathbf{u}}-\mu\mathbf{J}_{\phi}+\partial_{\nabla\phi}f\partial_{t}^{\bullet}\phi\big{]}\\ &\quad+\big{[}\partial_{\phi}f-\textnormal{div}\,(\partial_{\nabla\phi}f)-\mu\big{]}\partial_{t}^{\bullet}\phi+\nabla\mu\cdot\mathbf{J}_{\phi}\\ &\quad-\big{[}\mathbf{T}+p\mathbf{I}+\nabla\phi\otimes\partial_{\nabla\phi}f\big{]}:\nabla{\mathbf{u}}&&\quad\text{in $Q_{T}$.}\end{aligned}$ (3.56)
@@
$\mathbf{T}&=\mathbf{S}-p\mathbf{I}-\varepsilon\nabla\phi\otimes\nabla\phi\\ &=2\nu(\phi)\mathbf{D}{\mathbf{u}}-p\mathbf{I}-\varepsilon\nabla\phi\otimes\nabla\phi\quad\text{in $Q_{T}$.}\end{aligned}$ (3.64)
@@ 3.3.2 Computations on the boundary
$0&\geq\partial_{\phi}g\,\partial_{t}^{\circ}\phi+\partial_{\psi}g\,\partial_{t}^{\circ}\psi+\partial_{\nabla_{\Gamma}\psi}g\cdot\partial_{t}^{\circ}\nabla_{\Gamma}\psi+\textnormal{div}_{\Gamma}\,\mathbf{K}_{e}+H\mathbf{K}_{e}\cdot{\mathbf{n}}\\ &\quad-\tfrac{1}{2}(\mathbf{J}\cdot{\mathbf{n}}){\mathbf{u}}\cdot{\mathbf{u}}-\mathbf{T}{\mathbf{n}}\cdot{\mathbf{u}}-\mu\mathbf{J}_{\phi}\cdot{\mathbf{n}}+\varepsilon\,\partial_{\mathbf{n}}\phi\,\partial_{t}^{\circ}\phi\\ &\quad-\theta\partial_{t}^{\circ}\psi-\textnormal{div}_{\Gamma}(\mathbf{K}_{\psi}\theta)+\nabla_{\Gamma}\theta\,\mathbf{K}_{\psi}+\theta\mathbf{J}_{\psi}\cdot{\mathbf{n}}\\ &\quad-\textnormal{div}_{\Gamma}(q{\mathbf{u}})+\nabla_{\Gamma}q\cdot{\mathbf{u}}\end{aligned}$ (3.68)
@@
[…]
@@
$&\partial_{\nabla_{\Gamma}\psi}g\cdot\partial_{t}^{\circ}\nabla_{\Gamma}\psi\\ &\quad=\partial_{\nabla_{\Gamma}\psi}g\,\nabla_{\Gamma}\partial_{t}^{\circ}\psi+\partial_{\nabla_{\Gamma}\psi}g\cdot\Big{\{}\nabla_{\Gamma}{\mathbf{u}}\nabla_{\Gamma}\psi-\mathbf{P}_{\Gamma}\big{[}\nabla_{\Gamma}{\mathbf{u}}+(\nabla_{\Gamma}{\mathbf{u}})^{T}\big{]}\nabla_{\Gamma}\psi\Big{\}}\\ &\quad=\partial_{\nabla_{\Gamma}\psi}g\,\nabla_{\Gamma}\partial_{t}^{\circ}\psi-\delta\big{(}\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi\big{)}:\nabla_{\Gamma}{\mathbf{u}}\\[4.30554pt] &\quad=\textnormal{div}_{\Gamma}\big{[}\partial_{\nabla_{\Gamma}\psi}g\partial_{t}^{\circ}\psi-\delta(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi){\mathbf{u}}\big{]}-\textnormal{div}_{\Gamma}\big{(}\partial_{\nabla_{\Gamma}\psi}g\big{)}\partial_{t}^{\circ}\psi+\delta\,\textnormal{div}_{\Gamma}(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi){\mathbf{u}}\end{aligned}$ (3.71)
@@
$\delta(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi){\mathbf{u}}\cdot{\mathbf{n}}=\nabla_{\Gamma}\Big{(}\frac{\delta}{2}\left|\nabla_{\Gamma}\psi\right|^{2}\Big{)}\cdot{\mathbf{n}}+\Delta_{\Gamma}\psi\nabla_{\Gamma}\psi\cdot{\mathbf{n}}=0$ (3.72)

### v2

@@ 2.3  The role of dynamic boundary conditions in our model
  • For $K=0$ and $L\in(0,\infty)$, the non-convective bulk-surface Cahn–Hilliard equation was proposed in [59] to interpolate between the models associated with $K=0$, $L=0$ and K=0{K=0}, $L=\infty$}. In this case, a transfer of material between bulk and surface will occur, and the number $L^{-1}$ is related to the rate of absorption (or adsorption) of bulk material by the boundary (cf. [59]).
@@ 2.4  Variants of our model with a static boundary
$\partial_{t}\big(\rho(\phi){\mathbf{u}}\big)+\textnormal{div}\,\big({\mathbf{u}}\otimes(\rho(\phi){\mathbf{u}}+\mathbf{J})\big)=\textnormal{div}\,(\mathbf{T}),\quad\textnormal{div}\,{\mathbf{u}}=0,$ $\quad\text{in }Q_{T},$ (2.4a)
$\partial_{t}\phi+\textnormal{div}\,(\phi{\mathbf{u}})=\textnormal{div}\,(m_{\Omega}(\phi)\nabla\mu),$ $\quad\text{in }Q_{T},$ (2.4b)
$\mu=-\varepsilon\Delta\phi+\varepsilon^{-1}F^{\prime}(\phi),$ $\quad\text{in }Q_{T},$ (2.4c)
$\partial_{t}\psi+\textnormal{div}_{\Gamma}(\psi{\mathbf{u}}_{\tau})=\textnormal{div}_{\Gamma}(m_{\Gamma}(\psi)\nabla_{\Gamma}\theta)-\beta m_{\Omega}(\phi)\partial_{\mathbf{n}}\mu,$ $\quad\text{on }S_{T},$ (2.4d)
$\theta=-\delta\Delta_{\Gamma}\psi+\delta^{-1}G^{\prime}(\psi)+\alpha\varepsilon\partial_{\mathbf{n}}\phi,$ $\quad\text{on }S_{T},$ (2.4e)
$K\varepsilon\partial_{\mathbf{n}}\phi=\alpha\psi-\phi,\quad Lm_{\Omega}(\phi)\partial_{\mathbf{n}}\mu=\beta\theta-\mu,\quad K,L\in[0,\infty],$ $\quad\text{on }S_{T},$ (2.4f)
$\big[\mathbf{S}{\mathbf{n}}+\nabla_{\Gamma}q+\gamma_{\tau}{\mathbf{u}}\big]_{\tau}=\Big[\theta\nabla_{\Gamma}\psi+\frac{1}{2}\,(\mathbf{J}\cdot{\mathbf{n}})\,{\mathbf{u}}\Big]_{\tau},$ $\quad\text{on }S_{T},$ (2.4g)
$\textnormal{div}_{\Gamma}{\mathbf{u}}=0,\quad{\mathbf{u}}\cdot{\mathbf{n}}=0,$ $\quad\text{on }S_{T},$ (2.4h)
@@ 3  Model derivation
where for any $(t,{\mathbf{x}})\in\overline{Q_{T}}$, the flow map $0,T]\ni s\mapsto\eta_{s}(t,{\mathbf{x}})\in\overline{\Omega(s)}$ is the unique solution of the ordinary differential equation
@@
${\mathbf{u}}\cdot{\mathbf{n}}=\mathcal{V}\quad\text{on $S_{T}$}.$ (3.1)
@@ 3.1  Considerations based on local mass balance laws
Here, $\widehat{\mathbf{J}}_{i}$ are the mass fluxes corresponding to the motion of the two materials in the bulk. In our model, we also want to account for a transfer of material between bulk and surface. This is, for instance, needed in order to describe absorption processes. Therefore, we also consider functions $\sigma_{i}:S_{T}\to\mathbb{R}$, $i=1,2$, which represent the mass densities of the substrates on the free boundary. In general, the materials on the surface might differ from those in the bulk. This is the case, for example, if the materials are transformed by chemical reactions occurring at the boundary. Therefore, we interpret the densities $\sigma_{i}$ as independent functions that are not necessarily related to the traces of the densities $\rho_{i}$ (up to a constant carrying physical units), respectively. We assume that the the following mass balance law holds on the free boundary:
@@
$\begin{aligned} \rho(\partial_{t}{\mathbf{u}}+{\mathbf{u}}\cdot\nabla{\mathbf{u}})&=\textnormal{div}\,\,\tilde{\mathbf{T}}+(\textnormal{div}\,\,\mathbf{J}){\mathbf{u}}=\textnormal{div}\,(\tilde{\mathbf{T}}+{\mathbf{u}}\otimes\mathbf{J})-\nabla{\mathbf{u}}\;\mathbf{J}\quad\text{in $Q_{T}$.}\end{aligned}$ (3.23)
@@
$\begin{aligned} \textnormal{div}\,{\mathbf{u}}=\textnormal{div}\,\left(\frac{\rho_{1}}{\tilde{\rho}_{1}}{\mathbf{u}}_{1}+\frac{\rho_{2}}{\tilde{\rho}_{2}}{\mathbf{u}}_{2}\right)=-\partial_{t}\left(\frac{\rho_{1}}{\tilde{\rho}_{1}}+\frac{\rho_{2}}{\tilde{\rho}_{2}}\right)=-\partial_{t}1=0&&\quad\text{in $Q_{T}$,}\end{aligned}$ (3.26)
@@
$\begin{aligned} &\textnormal{div}_{\Gamma}\,{\mathbf{u}}=\textnormal{div}_{\Gamma}\left(\frac{\sigma_{1}}{\tilde{\sigma}_{1}}{\mathbf{u}}_{1}+\frac{\sigma_{2}}{\tilde{\sigma}_{2}}{\mathbf{u}}_{2}\right)\\ &\quad=-\partial_{t}^{\circ}\left(\frac{\sigma_{1}}{\tilde{\sigma}_{1}}+\frac{\sigma_{2}}{\tilde{\sigma}_{2}}\right)+{\mathbf{u}}_{\tau}\cdot\nabla_{\Gamma}\left(\frac{\sigma_{1}}{\tilde{\sigma}_{1}}+\frac{\sigma_{2}}{\tilde{\sigma}_{2}}\right)+\left(\frac{\widehat{\mathbf{J}}_{\Gamma,1}}{\tilde{\sigma}_{1}}+\frac{\widehat{\mathbf{J}}_{\Gamma,2}}{\tilde{\sigma}_{2}}\right)\cdot{\mathbf{n}}\\ &\quad=\left(\frac{\widehat{\mathbf{J}}_{\Gamma,1}}{\tilde{\sigma}_{1}}+\frac{\widehat{\mathbf{J}}_{\Gamma,2}}{\tilde{\sigma}_{2}}\right)\cdot{\mathbf{n}}=0&&\quad\text{on $S_{T}$.}\end{aligned}$ (3.27)
@@
$\tilde{\mathbf{J}}_{\Gamma,i}$ $=\frac{\mathbf{J}_{\Gamma,i}}{\tilde{\sigma}_{i}},$ $\quad i=1,2,$ $\qquad\mathbf{J}_{\psi}$ $=\tilde{\mathbf{J}}_{\Gamma,2}-\tilde{\mathbf{J}}_{\Gamma,1}$ on $S_{T}$, (3.29)
@@
$\begin{aligned} \textnormal{div}\,\,\mathbf{J}&=\frac{1}{2}\bigg[\tilde{\rho}_{2}\textnormal{div}\,\,\tilde{\mathbf{J}}_{2}+\tilde{\rho}_{1}\textnormal{div}\,\,\tilde{\mathbf{J}}_{1}\bigg]+\frac{1}{2}\bigg[\tilde{\rho}_{2}\textnormal{div}\,\,\tilde{\mathbf{J}}_{2}+\tilde{\rho}_{1}\textnormal{div}\,\,\tilde{\mathbf{J}}_{1}\bigg]\\[4.30554pt] &=\frac{\tilde{\rho}_{2}-\tilde{\rho}_{1}}{2}\textnormal{div}\,\,\tilde{\mathbf{J}}_{2}-\frac{\tilde{\rho}_{2}-\tilde{\rho}_{1}}{2}\textnormal{div}\,\,\tilde{\mathbf{J}}_{1}=\frac{\tilde{\rho}_{2}-\tilde{\rho}_{1}}{2}\textnormal{div}\,\,\mathbf{J}_{\phi}&&\quad\text{in $Q_{T}$.}\end{aligned}$ (3.35)
@@
Assuming that $\phi$ and $\psi$ only attain values in the interval [1,1][-1,1], we use (3.9), (3.10) and (3.17) to conclude that the density $\rho=\rho(\phi)$ is given by
@@ 3.2  Local energy dissipation laws
$\begin{aligned} 0&\geq\frac{\mathrm{d}}{\mathrm{d}t}\left[\int_{V(t)}e_{\Omega}({\mathbf{u}},\phi,\nabla\phi)\;\mathrm{d}{\mathbf{x}}+\int_{\partial V(t)\cap\Gamma(t)}e_{\Gamma}(\phi,\psi,\nabla_{\Gamma}\psi)\;\mathrm{d}\mathcal{H}^{d-1}\right]\\ &\qquad+\int_{\partial V(t)\cap\Omega(t)}\mathbf{J}_{e}\cdot{\mathbf{n}}_{\partial V(t)}\;\mathrm{d}\mathcal{H}^{d-1}+\int_{\partial(\partial V(t)\cap\Gamma(t))}\mathbf{K}_{e}\cdot{\mathbf{n}}_{\partial(\partial V(t)\cap\Gamma(t))}\;\mathrm{d}\mathcal{H}^{d-2}.\end{aligned}$ (3.43)
@@
$\begin{aligned} 0&\geq\frac{\mathrm{d}}{\mathrm{d}t}\left[\int_{V(t)}e_{\Omega}({\mathbf{u}},\phi,\nabla\phi)\;\mathrm{d}{\mathbf{x}}+\int_{\partial V(t)\cap\Gamma(t)}e_{\Gamma}(\phi,\psi,\nabla_{\Gamma}\psi)\;\mathrm{d}\mathcal{H}^{d-1}\right]\\ &\qquad+\int_{V(t)}\textnormal{div}\,\,\mathbf{J}_{e}\;\mathrm{d}{\mathbf{x}}+\int_{\partial V(t)\cap\Gamma(t)}\textnormal{div}_{\Gamma}\,\mathbf{K}_{e}+H\mathbf{K}_{e}\cdot{\mathbf{n}}-\mathbf{J}_{e}\cdot{\mathbf{n}}\;\mathrm{d}\mathcal{H}^{d-1}\,.\end{aligned}$ (3.44)
@@
$\begin{aligned} 0&\geq\int_{V(t)}\partial_{t}e_{\Omega}({\mathbf{u}},\phi,\nabla\phi)+\textnormal{div}\,\big(e_{\Omega}({\mathbf{u}},\phi,\nabla\phi){\mathbf{u}}\big)+\textnormal{div}\,\,\mathbf{J}_{e}\;\mathrm{d}{\mathbf{x}}\\ &\quad+\int_{\partial V(t)\cap\Gamma(t)}\partial_{t}^{\circ}e_{\Gamma}(\phi,\psi,\nabla_{\Gamma}\psi)\;\mathrm{d}\mathcal{H}^{d-1}\\ &\quad+\int_{\partial V(t)\cap\Gamma(t)}\textnormal{div}_{\Gamma}\,\mathbf{K}_{e}+H\mathbf{K}_{e}\cdot{\mathbf{n}}-\mathbf{J}_{e}\cdot{\mathbf{n}}\;\mathrm{d}\mathcal{H}^{d-1}.\end{aligned}$ (3.47)
@@ 3.3.1 Computations in the bulk
$\begin{aligned} 0&\geq\textnormal{div}\,\big[\mathbf{J}_{e}-\tfrac{1}{2}\left|{\mathbf{u}}\right|^{2}\mathbf{J}+\mathbf{T}^{T}{\mathbf{u}}-\mu\mathbf{J}_{\phi}+\partial_{\nabla\phi}f\partial_{t}^{\bullet}\phi\big]\\ &\quad+\big[\partial_{\phi}f-\textnormal{div}\,(\partial_{\nabla\phi}f)-\mu\big]\partial_{t}^{\bullet}\phi+\nabla\mu\cdot\mathbf{J}_{\phi}\\ &\quad-\big[\mathbf{T}+p\mathbf{I}+\nabla\phi\otimes\partial_{\nabla\phi}f\big]:\nabla{\mathbf{u}}&&\quad\text{in $Q_{T}$.}\end{aligned}$ (3.56)
@@
$\begin{aligned} \mathbf{T}&=\mathbf{S}-p\mathbf{I}-\varepsilon\nabla\phi\otimes\nabla\phi\\ &=2\nu(\phi)\mathbf{D}{\mathbf{u}}-p\mathbf{I}-\varepsilon\nabla\phi\otimes\nabla\phi\quad\text{in $Q_{T}$.}\end{aligned}$ (3.64)
@@ 3.3.2 Computations on the boundary
$\begin{aligned} 0&\geq\partial_{\phi}g\,\partial_{t}^{\circ}\phi+\partial_{\psi}g\,\partial_{t}^{\circ}\psi+\partial_{\nabla_{\Gamma}\psi}g\cdot\partial_{t}^{\circ}\nabla_{\Gamma}\psi+\textnormal{div}_{\Gamma}\,\mathbf{K}_{e}+H\mathbf{K}_{e}\cdot{\mathbf{n}}\\ &\quad-\tfrac{1}{2}(\mathbf{J}\cdot{\mathbf{n}}){\mathbf{u}}\cdot{\mathbf{u}}-\mathbf{T}{\mathbf{n}}\cdot{\mathbf{u}}-\mu\mathbf{J}_{\phi}\cdot{\mathbf{n}}+\varepsilon\,\partial_{\mathbf{n}}\phi\,\partial_{t}^{\circ}\phi\\ &\quad-\theta\partial_{t}^{\circ}\psi-\textnormal{div}_{\Gamma}(\mathbf{K}_{\psi}\theta)+\nabla_{\Gamma}\theta\,\mathbf{K}_{\psi}+\theta\mathbf{J}_{\psi}\cdot{\mathbf{n}}\\ &\quad-\textnormal{div}_{\Gamma}(q{\mathbf{u}})+\nabla_{\Gamma}q\cdot{\mathbf{u}}\end{aligned}$ (3.68)
@@
$\begin{aligned} &\partial_{\nabla_{\Gamma}\psi}g\cdot\Big\{\nabla_{\Gamma}{\mathbf{u}}\nabla_{\Gamma}\psi-\mathbf{P}_{\Gamma}\big[\nabla_{\Gamma}{\mathbf{u}}+(\nabla_{\Gamma}{\mathbf{u}})^{T}\big]\nabla_{\Gamma}\psi\Big\}\\ &\quad=-\delta\nabla_{\Gamma}\psi\cdot(\nabla_{\Gamma}{\mathbf{u}})^{T}\nabla_{\Gamma}\psi=-\delta\big(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi\big):\nabla_{\Gamma}{\mathbf{u}}\end{aligned}$ (3.70)
@@
$\begin{aligned} &\partial_{\nabla_{\Gamma}\psi}g\cdot\partial_{t}^{\circ}\nabla_{\Gamma}\psi\\ &\quad=\partial_{\nabla_{\Gamma}\psi}g\,\nabla_{\Gamma}\partial_{t}^{\circ}\psi+\partial_{\nabla_{\Gamma}\psi}g\cdot\Big\{\nabla_{\Gamma}{\mathbf{u}}\nabla_{\Gamma}\psi-\mathbf{P}_{\Gamma}\big[\nabla_{\Gamma}{\mathbf{u}}+(\nabla_{\Gamma}{\mathbf{u}})^{T}\big]\nabla_{\Gamma}\psi\Big\}\\ &\quad=\partial_{\nabla_{\Gamma}\psi}g\,\nabla_{\Gamma}\partial_{t}^{\circ}\psi-\delta\big(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi\big):\nabla_{\Gamma}{\mathbf{u}}\\[4.30554pt] &\quad=\textnormal{div}_{\Gamma}\big[\partial_{\nabla_{\Gamma}\psi}g\partial_{t}^{\circ}\psi-\delta(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi){\mathbf{u}}\big]-\textnormal{div}_{\Gamma}\big(\partial_{\nabla_{\Gamma}\psi}g\big)\partial_{t}^{\circ}\psi+\delta\,\textnormal{div}_{\Gamma}(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi){\mathbf{u}}\end{aligned}$ (3.71)
@@
$\delta(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi){\mathbf{u}}\cdot{\mathbf{n}}=0$ (3.72)